# Supervised Learning-enhanced Multi-Group Actor Critic for Live Stream Allocation in Feed

### v5

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2, August 3–7, 2025, Toronto, ON, Canada; Toronto, ON; Canadaisbn: 979-8-4007-1245-6/25/08ccs: Information systems Recommender systemsccs: Computing methodologies Reinforcement learning
@@ 1. Introduction
However, the intrinsic issues of divergence and instability associated with traditional reinforcement learning algorithms (RL) (François-Lavet et al., 2015; Fujimoto et al., 2018; Kumar et al., 2019; Dasagi et al., 2019) are significantly exacerbated in the short video & live stream mixed recommendation system. A primary possible reason for this is the drastic fluctuations in both the live stream supply scale and user interaction behaviors over time, as shown in Fig. 7 in Appendix B. RL models often struggle to learn effective policies from data that exhibit such high variance across both time scales and user scales. In practice, we observe that RL models frequently encounter issues of policy deterioration or model collapse (Dohare et al., 2023). Most importantly, the live-stream allocation agent is the final module of the live-stream RS, as shown in Fig 1. If the RL agent is very instable or collapse in the online environment, we may inject too many live-streams in the short-video feed, which will cause severe malfunctions and will highly affect the user retention. Hence, in the industrial RS, this RL application is more risky than existing RL methods (Cai et al., 2023; Zhang et al., 2022, 2024), which are only applied in the multi-rank score aggregation of the ranking stage.
@@
To address the aforementioned issues, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm, SL-MGAC. Given the significant variation in users’ interest in live stream content, which introduces high variance at both the feature and feedback levels, we incorporate separate multi-group state decomposition (MG-SD) modules within both the actor and critic networks. Additionally, we combine multi-task supervised reward learning with traditional critic learning to not only restrict temporal difference (TD) error accumulation during model training but also improve the accuracy of Q-value estimation. Through experiments, we compare the SL-MGAC method with competitive baselines using offline policy evaluation (OPE) (Uehara et al., 2022) and online A/B tests to demonstrate its effectiveness and stability.
@@ 3.2. Multi-Group State Decomposition Module
Empirically, we show the effectiveness of this module in enhancing the stability of the RL model in subsequent experiments. We find that SL-MGAC without the MG-SD module tends to be more unstable during training, with a larger variance in its Q-values.
@@ 4.3. Implementation Details
To ensure fairness among all compared methods, we use a consistent feature extraction module with a 5000-size embedding layer. In addition to the embedding layer, the feature extraction module includes a 2-layer MLP with hidden sizes of [256, 128]. The batch size and number of epochs for all methods are 2048 and 500, respectively. The learning rate for the embedding layer is set to 1e-5, while the learning rate for other hidden parameters is 1e-3. The discount factor $\gamma$ is set to 0.9 for all methods and $\lambda$ in Eq 5 is 0.1. Detailed network architecture and parameter settings for SL-MGAC are provided in Table 4 of the Appendix A.
@@ 4.5. Ablation Study
Next, we compare the training processes of SL-MGAC and SL-MGAC (w/o MG), as shown in Fig. 4. We observe that the Q-value curve of SL-MGAC is more stable than that of SL-MGAC (w/o MG), and the Q-value variance of SL-MGAC is much smaller. This demonstrates the effectiveness and variance reduction effect of the MG-SD module.
Figure 4. The Q value curves between SL-MGAC and SL-MGAC (w/o MG) over 10 rounds of training. The lines correspond to the means of Q-value and the shaded areas correspond to the standard deviations.
@@ 4.7. Online A/B Test Experiment
We compare the improvements of different methods in terms of daily active live stream users (DAU), live stream watch time, video watch time, app usage duration, and user retention relative to the baseline. The baseline uses the SAC framework for live stream allocation. The results of the online A / B test are shown in Table 3.
@@ 4.8. Online Model Stability
We evaluate the online model stability between the SAC-based baseline and SL-MGAC. As shown in Fig. 6, the live stream allocation ratio fluctuates drastically throughout the day for both models. However, SL-MGAC exhibits greater stability than the baseline, with a smaller amplitude. Note that the live stream allocation ratio refers to the proportion of requests with action $a=1$.

### v6

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737264isbn: 979-8-4007-1454-2/2025/08ccs: Information systems Recommender systemsccs: Computing methodologies Reinforcement learning
@@ 1. Introduction
However, the intrinsic issues of divergence and instability associated with traditional reinforcement learning algorithms (RL) (François-Lavet et al., 2015; Fujimoto et al., 2018; Kumar et al., 2019; Dasagi et al., 2019) are significantly exacerbated in the short video & live stream mixed recommendation system. A primary possible reason for this is the drastic fluctuations in both the live stream supply scale and user interaction behaviors over time, as shown in Fig. 7 in Appendix B. RL models often struggle to learn effective policies from data that exhibit such high variance across both time scales and user scales. In practice, we observe that RL models frequently encounter issues of policy deterioration or model collapse (Dohare et al., 2023). Most importantly, the live-stream allocation agent is the final module of the live-stream RS, as shown in Fig 1. If the RL agent becomes highly unstable or collapses in the online environment, it may lead to an excessive injection of live-stream content into the short-video feed. In turn, this can negatively impact the exposure of advertisements and E-commerce videos, potentially causing severe system malfunctions and significant platform losses. Hence, in the industrial RS, this RL application is more risky than existing RL methods (Cai et al., 2023; Zhang et al., 2022, 2024), which are only applied in the multi-rank score aggregation of the ranking stage.
@@
To address the aforementioned problems, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm, SL-MGAC. Given the significant variation in users’ interest in live stream content, which introduces high variance at both the feature and feedback levels, we incorporate separate Multi-Group State Decomposition (MG-SD) modules within both the actor and critic networks. Additionally, we combine multi-task supervised reward learning with traditional critic learning to not only restrict temporal difference (TD) error accumulation during model training but also improve the accuracy of Q-value estimation. Through experiments, we compare the SL-MGAC method with competitive baselines using offline policy evaluation (OPE) (Uehara et al., 2022) and online A/B tests to demonstrate its effectiveness and stability.
@@ 3.2. Multi-Group State Decomposition Module
Empirically, we show the effectiveness of this module in enhancing the stability of the RL model in subsequent experiments. We find that SL-MGAC without the MG-SD module tends to be more unstable during training, with a larger variance in its Q-values. Furthermore, the MG-SD module is flexible and transferable, allowing the user-group partitioning strategy to be easily adapted to various scenarios in advertising and E-commerce.
@@ 3.4. Online Exploration and Deployment
Most importantly, we focus on reducing computational costs for successful deployment in two key aspects. First, for real-time training, we avoid using user-item cross features, which are commonly employed in industrial ranking models, thereby reducing the sparse embedding parameters. Additionally, the network parameters of the Actor part in SL-MGAC are much smaller than those of the Critic, with the complex Critic acting as a teacher to guide the Actor’s learning, as explained in Section 3.3.3. As a result, the total parameters of SL-MGAC are under 100 million in our recommendation system, significantly smaller than the parameter scale of typical ranking models, which usually exceed 1 billion parameters.
For fast online inference, the smaller Actor network enables fast inference, processing each request in under 20ms, even with over 10,000 requests per second during peak times. Furthermore, since the complex Critic is not used during online inference, this further reduces computational costs, contributing to faster response time in the online environment.
@@ 4.2.1. Baselines
  • RLUR (Cai et al., 2023). An RL model that aims to optimize the long-term user retention.
@@ 4.3. Implementation Details
To ensure fairness among all compared methods, we use a consistent feature extraction module with a 5000-size embedding layer. In addition to the embedding layer, the feature extraction module includes a 2-layer MLP with hidden sizes of [256, 128]. The batch size and number of epochs for all methods are 2048 and 500, respectively. The learning rate for the embedding layer is set to 1e-5, while the learning rate for other hidden parameters is 1e-3. The discount factor $\gamma$ is set to 0.9 for all methods and $\lambda$ in Eq 5 is 0.1. Detailed network architecture and parameter settings for SL-MGAC are provided in Table 4 of the Appendix A. And the code of SL-MGAC method is available at https://github.com/frankg1/SL-MGAC-torch.
@@ 4.5. Ablation Study
Next, we compare the training processes of SL-MGAC and SL-MGAC (w/o MG), as shown in Fig. 4. Note that the shaded area in yellow corresponds to the Q value std range of the SL-MGAC model, while the shaded area in purple corresponds to the Q value std range of the SL-MGAC model(w/o MG). We observe that the Q-value curve of SL-MGAC is more stable than that of SL-MGAC (w/o MG), and the Q-value variance of SL-MGAC is much smaller. This demonstrates the effectiveness and variance reduction effect of the MG-SD module.
Figure 4. The Q value curves between SL-MGAC and SL-MGAC (w/o MG) over 10 rounds of training. The lines correspond to the means of Q-value and the shaded areas correspond to the standard deviations (std).
@@ 4.7. Online A/B Test Experiment
The online A/B Test is conducted over a 5-day period in 2024 September and we randomly choose 20% users (over 20 million) in Kwai app as the experimental group. Then we compare the improvements of different methods in terms of daily active live stream users (DAU), live stream watch time, video watch time, app usage duration, and user retention relative to the baseline. The baseline uses the SAC framework for live stream allocation. The results of the online A/B test are shown in Table 3.
@@ 4.8. Online Model Stability
We evaluate the online model stability between the SAC-based baseline and SL-MGAC. The live-stream injection ratio is chosen as the evaluation metric, calculated by dividing the number of live-stream injection requests by the total number of requests within a given timestamp. As shown in Fig. 6, the live-stream allocation ratio for both models fluctuates significantly throughout the day, from high-traffic hours (such as 6:00–11:00) to low-traffic hours (14:00–18:00). This variation is due to changes in the overall user scale during these periods, which in turn affect the data distribution. However, SL-MGAC exhibits greater stability than the baseline, with a smaller amplitude. Note that the live stream allocation ratio refers to the proportion of requests with action $a=1$.
@@ 5.1. RL in Recommendation Systems
However, the aforementioned RL methods may encounter instability issues and could fail when faced with highly noisy or drastically changing data distributions. The proposed SL-MGAC method is more robust and has been successfully deployed in the ”risky” final stage of a real-world industrial RS. We believe this work provides valuable insights not only into the model framework design for RL applications in industrial RS, but also offers practical experience in deploying robust and successful RL systems for other industrial companies.
@@ 6. Conclusion
7. Acknowledgment
The authors acknowledge Wei Bai, Xiaoshuang Chen and anonymous reviewers for proposing detailed modification advice to help us improve the quality of this manuscript.